OpenStudy (anonymous):

Look at the right triangle ABC. A student made the following chart to prove that AB2 + BC2 = Statement Justificat 1. Triangle ABC is similar to triangle BDC 1. Angle ABC = Angle BCD and Angle BCA = Angle DBC 2. BC2 = AC × DC 2. BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC 3. Triangle ABC is similar to triangle ABD 3. Angle ABC = Angle ADB and Angle BAC = Angle BAD 4. AB2 = AC × AD 4. AB ÷ AD = AC ÷ AB because triangle ABC is similar to triangle ABD 5. AB2 + BC2 = AC × AD + AC × DC = AC (AD + DC) 5. Adding Statement

OpenStudy (anonymous):

Statement Justification 1. Triangle ABC is similar to triangle BDC 1. Angle ABC = Angle BCD and Angle BCA = Angle DBC 2. BC2 = AC × DC 2. BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC 3. Triangle ABC is similar to triangle ABD 3. Angle ABC = Angle ADB and Angle BAC = Angle BAD 4. AB2 = AC × AD 4. AB ÷ AD = AC ÷ AB because triangle ABC is similar to triangle ABD 5. AB2 + BC2 = AC × AD + AC × DC = AC (AD + DC) 5. Adding Statement 1 and Statement 2 6. AB2 + BC2 = AC2 6. AD + DC = AC What is the flaw in the student’s proof? Justification 4 should be “AB ÷ AD = AB ÷ AC because triangle ABC is similar to triangle ABD”. Justification 1 should be “Angle ABC = Angle BDC and Angle BCA = Angle BCD”. Justification 2 should be “BC ÷ DC = BC ÷ AC because triangle ABC is similar to triangle BDC”. Justification 3 should be “Angle ABC = Angle BAD and Angle BAC = Angle ABD”.